# Preprint, to appear in Journal for Geometry and Graphics, Vol. 29/2

### v2

@@ Abstract.
We present a simplified proof of a forty-year-old result concerning the tiling of the plane with equilateral convex polygons. Our approach is based on a theorem by M. Rao, who used an exhaustive computer search to confirm the completeness of the well-known list of fifteen pentagon types. Assuming the validity of Rao’s result, we provide a concise and mainly geometrical proof of a tiling theorem originally due to Hirschhorn and Hunt. Finally, a possible connection to quasicrystals is sketched.
@@ 1. Introduction
Assuming that Rao’s classification is indeed complete, we can revisit the original problem from a new angle: With the list of fifteen pentagon types accepted as exhaustive, we propose a significantly simpler and mainly geometric approach for characterizing all equilateral convex polygons that tile the plane. The main result of this paper is a geometrical proof of the following proposition:
@@ 2. Proofs of lemmas and proposition
From $D+C/2=\pi=(D-\pi/2)+C/2+\pi/2$ it follows that $B$ must lie on a line through $D$ perpendicular to $e$; $\Rightarrow A$ cannot lie lower than vertex $E;\;\;\Rightarrow$ angle $E\leq\pi/2;\Rightarrow$ with $E+B/2=\pi$ it follows that $B/2\geq\pi/2;\;\;\Rightarrow$ due to convexity $B/2=\pi/2;\;\;\Rightarrow E=\pi/2$, satisfied only by the solution with dotted lines;    $\Rightarrow E+B/2<\pi$ which produces a contradiction. ∎
@@ 3.2. Type 6
We assume that a convex case exists with all side lengths equal to 1. $\sin(C)$ can be expressed by $2\,\sin(C/2)\cos(C/2)$ and with the substitution $z=\sin^{2}(C/2)$ we end up with the quadratic equation $z^{2}-3z/4-1/16=0$ which produces only one positive solution $z=(3+\sqrt{13})/8$ and $C\!\approx 130.65^{\circ}$. Hence, $A=2C>2\pi$, which is a contradiction to the assumption.

### v3

# Preprint, to appear in Journal for Geometry and Graphics, Vol. 29 (2)

@@ Abstract.
We present a simplified proof of a forty-year-old result concerning the tiling of the plane with equilateral convex polygons. Our approach is based on a theorem by M. Rao, who used an exhaustive computer search to confirm the completeness of the well-known list of fifteen pentagon types. Assuming the validity of Rao’s result, we provide a concise and mainly geometric proof of a tiling theorem originally due to Hirschhorn and Hunt. Finally, a possible connection to quasicrystals is sketched.
@@ 1. Introduction
Assuming that Rao’s classification is indeed complete, we can revisit the original problem from a new angle: With the list of fifteen pentagon types accepted as exhaustive, we propose a significantly simpler and mainly geometric approach for characterizing all equilateral convex polygons that tile the plane. The main result of this paper is a geometric proof of the following proposition:
@@ 2. Proofs of lemmas and proposition
From $D=\pi-C/2$ it follows that $B$ must lie on a line through $D$ perpendicular to $e$; $\Rightarrow A$ cannot lie lower than vertex $E$, since $|b|=|e|;\;\;\Rightarrow$ angle $E\leq\pi/2;\Rightarrow$ with $E+B/2=\pi$ it follows that $B/2\geq\pi/2;\;\;\Rightarrow$ due to convexity $B/2=\pi/2;\;\;\Rightarrow E=\pi/2$, satisfied only by the solution with dotted lines;    $\Rightarrow E+B/2<\pi$ which produces a contradiction. ∎
@@ 3.2. Type 6
We assume that a convex case exists with all side lengths equal to 1. $\sin(C)$ can be expressed by $2\,\sin(C/2)\cos(C/2)$ and with the substitution $z=\sin^{2}(C/2)$ we end up with the quadratic equation $z^{2}-3z/4-1/16=0$ which produces only one positive solution $z=(3+\sqrt{13})/8$ and $C\!\approx 130.65^{\circ}$. Hence, $A=2C>\pi$, which is a contradiction to the assumption.